Problem 1: Find a particular solution to the following equations +y= t + sint + 1let + Již + y = 2t sint dy-4d4 = e2t + + +5+Find a particular solution to the following equations
Answer #1

사2 The given differential equation is dy + y = ť+ sinttllet is. (03+) y = ²+ sinttllet Particular solution Jp = 1 (+²+ sint i& The given differential equation is de 2 + y = 2t sint ie. (D²+1)y=2tsint Particular solution: Yp = a (atsint) = 12 (2t & im③ The given differential equation is e - 4 edt +++54+4 ic. (05-403) y = (2+ + 8 + 5++4 Particular solution: Joosenz (eze+t+st